OpenStudy (anonymous):

solve y-2x=8 2y+2x=40

undeadknight26 (undeadknight26):

Is this all one question?

undeadknight26 (undeadknight26):

if y-2x=8 is a seperate question it would be y = 2x + 8

OpenStudy (anonymous):

@undeadknight26 Can you help me?

OpenStudy (anonymous):

The domain of the following relation: R: {(3, -2), (1, 2), (-1, -4), (-1, 2)} is {-1, 1, 3} {-1, -1, 1, 3} {-4, -2, 2, 2} {-4, -2, 2}

undeadknight26 (undeadknight26):

-1 1 3. first one...remember the domain is the x and when you are listening them you do not double list.

OpenStudy (anonymous):

The range of the following relation: R: {(3, -2), (1, 2), (-1, -4), (-1, 2)} is {-1, 1, 3} {-1, -1, 1, 3} {-4, -2, 2, 2} {-4, -2, 2}

OpenStudy (anonymous):

How about this one? @undeadknight26

undeadknight26 (undeadknight26):

YOu try this one i said that to domain is the first letter than what is the range?

OpenStudy (anonymous):

x is the first letter

undeadknight26 (undeadknight26):

yes as i said before the x and the domain is the same.

OpenStudy (anonymous):

Is it the second one?

undeadknight26 (undeadknight26):

please start up another question if you have more questions or email me.

OpenStudy (anonymous):

it just says solve using subsitution method y-2x=8 2y+2x=40

undeadknight26 (undeadknight26):

then they are separate questions.

OpenStudy (anonymous):

so how do i solve it

undeadknight26 (undeadknight26):

You are solving for y right?

OpenStudy (anonymous):

yes

undeadknight26 (undeadknight26):

0037109 i cannot email to you because it is not a valid username...please start your own question.

undeadknight26 (undeadknight26):

So we can put it into: y = 2x + 8 /2 /2 y = (x + 8) i believe this is how you answer it.

undeadknight26 (undeadknight26):

Are we still solving for y for the second one?

OpenStudy (anonymous):

yes

undeadknight26 (undeadknight26):

well then i am assuming: 2y = 2x - 40 /2 /2 2y = x - 20 /2 /2 y = x - 10? let me double check.

undeadknight26 (undeadknight26):

No apparently the answer is y = 20 -x

undeadknight26 (undeadknight26):

I have no idea how to do that srry lol...unless 2y = 2x - 40 /2 /2 /2 y = x - 20...

undeadknight26 (undeadknight26):

I hope i helped.